# Proactive Recommendation in Social Networks: Steering User Interest via Neighbor Influence

### v1

@@ 3.1. Identification
(3) $Y_{ui}\left(O_{ui}=1,\bm{O}_{\mathcal{N}_{ui}}=\bm{o}_{\mathcal{N}_{ui}}\right)=Y_{ui}\left(O_{ui}=1,\bm{G}_{ui}=\bm{g}_{ui}\right).$
Assumption 2.
@@
(4) $Y_{ui}\left(O_{ui}=1,\bm{G}_{ui}=\bm{g}_{ui}\right)\perp\left(O_{ui},\bm{G}_{ui}\right)\mid\bm{X}_{ui}.$
Assumption 3.
@@
[…]
Equations (a), (b), and (c) are based on Assumptions 1, 2, and 3, respectively. It is worth noting that the three standard assumptions commonly employed in causal inference are equally applicable in recommendation. Assumption 1 benefits significantly from the strong representational capabilities of graph convolution in recommendation (He et al., 2020), which can be adapted to capture information regarding the neighbors’ features and the exposure of the item to the neighbors. Assumptions 2 and 3 are adaptations of conventional assumptions in the traditional causal recommendation(Wu et al., 2022; Li et al., 2023c), tailored to account for the presence of interference.
3.2. Interference Representation-based
Estimation Module
@@ 3.2. Interference Representation-based
(7) $\bm{e}_{ui}=\bm{e}_{u}\odot\bm{e}_{i},$
where $\bm{e}_{u}$ and $\bm{e}_{i}$ are learnable embeddings for user $u$ and item $i$, respectively, which can be modeled by previous recommender models such as Matrix Factorization (MF) (Koren et al., 2009) and LightGCN (He et al., 2020).
In social networks, when $u$’s neighbor $u^{\prime}$ is exposed to item $i$, $u^{\prime}$ may be inclined to share their interests $\bm{e}_{u^{\prime}i}$ with $u$. Moreover, the social closeness between different neighbors and a user is not uniform resulting in differing degrees of influence on the user’s interests in the item. Therefore, the interference representation-based estimation module aggregates unit $(u,i)$’s neighbors’ treatment values $\bm{o}_{\mathcal{N}_{ui}}$ and user $u$’s neighbors’ individual interests in item $i$ as follows:
(8) $\bm{g}_{ui}=\bm{\mathrm{W}}\sum_{u^{\prime}\in\mathcal{N}_{u}}o_{u^{\prime}i}a_{uu^{\prime}}\bm{e}_{u^{\prime}i},$
@@
(9) $a_{uu^{\prime}}=\frac{\bm{w}^{\intercal}[\bm{e}_{u}\parallel\bm{e}_{u^{\prime}}]}{\sum_{u^{\prime\prime}\in\mathcal{N}_{u}}\bm{w}^{\intercal}[\bm{e}_{u}\parallel\bm{e}_{u^{\prime\prime}}]},$
where $\bm{w}$ is a learnable weight vector and [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] denotes the concatenation of two vectors.
Concatenating unit $(u,i)$’s individual representation and interference representation as $[\bm{e}_{ui}\parallel\bm{g}_{ui}]$, we can capture user’s $u$’s final inclination towards interacting with the item $i$. Any classifier can be used to model the interference potential outcome of unit $(u,i)$ (e.g., a linear classifier). In this work, we follow MF and model $P(Y_{ui}(O_{ui}=1,\bm{O}_{\mathcal{N}_{u}i}=\bm{o}_{\mathcal{N}_{u}i})=1\mid\bm{X}_{ui}=\bm{x}_{ui})$ as:
(10) $\sigma\left(\bm{1}^{\intercal}[\bm{e}_{ui}\parallel\bm{g}_{ui}]\right),$
where $\bm{1}$ is a vector where each element’s value is 1 and $\sigma(\cdot)$ is the sigmoid function. In estimating interference-aware potential outcome, we can specify item $i$’s exposure $\bm{o}_{\mathcal{N}_{u}i}$ to user $u$’s neighbors by observed data $\mathcal{D}$. The modeled interference-aware potential outcomes is denoted as $\hat{p}_{ui}$. We optimize the above learnable parameters by minimizing the following loss:
(11) $\frac{1}{|\mathcal{D}|}\sum_{(\bm{x}_{ui},o_{ui},y_{ui})\in\mathcal{D}}l(\hat{p}_{ui},y_{ui}),$
@@
3.3. Post-Learning-based Optimization Module
We propose a post-learning-based optimization module to achieve the second point in Section 2.2, which consists of: (1) estimating the sum of probabilities of positive potential outcomes of target units under interference; (2) designing a cost function to measure the damage to the neighbors resulting from the target item’s exposure; and (3) optimizing the treatment values of the neighbors to trade off steering performance and the damage to the neighbors.
3.3.1. Estimation for target units
@@
3.3.2. Design of cost function
We utilize an existing recommender model as the user experience model [^1] In practice, we train a LightGCN (He et al., 2020), a representative recommender model, as a user experience model, since we currently lack access to the recommendation platforms’ proprietary recommender models designed for user experience. and based on this, we design a cost function. The cost function measures the damage to a neighbor by measuring the decrease in the sum of ratings for items in the adjusted recommendation list from NIRec compared to the sum of ratings in the original recommendation list from the user experience model, both ratings as assigned by the user experience model. In particular, given the user experience model, we obtain: (1) an item list with the top k highest ratings to neighbor $u^{\prime}$; and (2) an exposure treatment value $o^{ex}_{u^{\prime}i^{\ast}}$ indicating that whether $i^{\ast}$ is in the top k list ($o^{ex}_{u^{\prime}i^{\ast}}=1$) or not ($o^{ex}_{u^{\prime}i^{\ast}}=0$). From the user experience model’s perspective, if we replace an item in the top k item list, it will damage $u^{\prime}$’s experience. Therefore, we design cost function $c(\cdot)$ as follows:
@@ 3.3.2. Design of cost function
3.3.3. Treatment values optimization
@@
(12) $\begin{split}&\sum_{u^{\ast}\in\mathcal{U^{\ast}}}P(Y_{u^{\ast}i^{\ast}}(\bm{O}_{\mathcal{U}^{\ast}i^{\ast}}=\bm{1},\bm{O}_{\mathcal{N}_{\mathcal{U}^{\ast}}i^{\ast}}=\bm{o}_{\mathcal{N}_{\mathcal{U}^{\ast}}i^{\ast}})=1\mid\bm{X}_{u^{\ast}i^{\ast}}=\bm{x}_{u^{\ast}i^{\ast}})\\ -&\gamma\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}c\left(o_{u^{\prime}i^{\ast}}\right).\end{split}$
@@
4. EXPERIMENT SETUP
Table 1. The Statistics of Ciao, Epinions, Filmtrust.
@@ 4. EXPERIMENT SETUP
#Influential User-Item Pairs 179,727 55,029 13,992
Since no offline datasets provide users’ feedback under varying neighbor interference, we propose a framework for user feedback simulation. It is noteworthy that semi-simulation is a viable and commonly practiced experimental method for recommendation tasks in the interactive evaluation settings (Sato et al., 2019; Gao et al., 2024; Sato et al., 2020; Gao et al., 2023; Zhu et al., 2023). Following the user feedback simulation, we present how to select target users and items for PRSN. Finally, we design evaluation metrics and baselines to reveal the effectiveness of NIRec for PRSN.
4.1. Real-World Datasets
We use two product review datasets, Ciao and Epinions (Tang et al., 2012a, b), as well as a movie review dataset Filmtrust (Guo et al., 2013), for simulating users’ feedback under varying neighbor interference, which results in three user feedback simulations. These datasets contain user IDs, item IDs, ratings (from 1 to 5), and a directed social network with trust relations. We use the 10-core setting for Ciao and Epinions and the 5-core setting [^2] The number of items in Filmtrust is on the order of 100, which is relatively small. for Filmtrust. The statistics of the datasets are shown in Table 1.
4.2. User Feedback Simulation
@@
4.2.1. Representation Simulation
We simulate representations for both users and items to generate user feedback under neighbor interference. To this end, we train an MF model using the observed ratings to predict and complete missing ratings in the user-item matrix. However, the model always falls into overestimation on missing ratings (Pan et al., 2023). To this end, we follow (Dai et al., 2022) to align the distribution of the ratings in the completed matrix with that of ratings in the random exposure dataset (e.g., Yahoo!R3 (Marlin et al., 2012)), thereby enhancing the reliability of our simulations. Thereafter, we use all ratings to train a new MF model to obtain our simulated user representation $\bm{\mathrm{E}}_{\mathcal{U}}^{\ast}$ and item representation $\bm{\mathrm{E}}_{\mathcal{I}}^{\ast}$.
4.2.2. User Feedback Generation Function
According to standard potential outcome simulation in causal inference (Ma and Tresp, 2021; Ma et al., 2022), we design a user feedback generation function considering neighbor interference as follows:
(13) $\begin{split}\sigma\left(\eta\left(\bm{e}_{u},\bm{e}_{i}\right)+\beta\Delta\left(\bm{\mathrm{E}}_{\mathcal{N}_{u}},\bm{e}_{u},\bm{e}_{i},\bm{o}_{\mathcal{N}_{u}i}\right)\right).\end{split}$
Here $\eta(\bm{e}_{u},\bm{e}_{i})$ is user $u$’s individual interests in interacting with item $i$ and shorted as $\eta_{ui}$ for brevity. $\Delta(\bm{\mathrm{E}}_{\mathcal{N}_{u}},\bm{e}_{u},\bm{e}_{i},\bm{o}_{\mathcal{N}_{ui}})$ is the neighbors’ interference effect and shorted as $\Delta_{ui}$. $\beta$ is the coefficient that controls the strength of the interference. In particular, we design $\eta_{ui}$ and $\Delta_{ui}$ under a linear setting:
(14) $\begin{split}\eta_{ui}&=\bm{e}_{u}\cdot\bm{e}_{i}-3,\\ \Delta_{ui}&=\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}}\sum_{u^{\prime}\in\mathcal{N}_{u}}{o_{u^{\prime}i}\sigma\left(\bm{e}_{u}\cdot\bm{e}_{u^{\prime}}\right)\eta_{u^{\prime}i}}.\end{split}$
Here, we consider a user to like an item if the corresponding rating is above 3; otherwise, the user is considered to dislike the item, given that the range of ratings fitted by MF for simulation is from 1 to 5. If the user’s neighbor’s rating of the item is higher than 3, the influence on the user is positive; otherwise, it is negative. Given $\bm{\mathrm{E}}_{\mathcal{U}}^{\ast}$ and $\bm{\mathrm{E}}_{\mathcal{I}}^{\ast}$, we set the positive potential outcome probability of unit $(u,i)$ as $\sigma\left(\eta\left(\bm{e}^{\ast}_{u},\bm{e}^{\ast}_{i}\right)+\beta\Delta\left(\bm{\mathrm{E}}^{\ast}_{\mathcal{N}_{u}},\bm{e}^{\ast}_{u},\bm{e}^{\ast}_{i},\bm{o}_{\mathcal{N}_{u}i}\right)\right)$.
4.2.3. Semi-Synthetic Datasets
We generate synthetic datasets with user feedback to train our interference representation-based estimation module as existing work (Dai et al., 2022). We adjust the propensity for observing the feedback to ensure that the generated data aligns as closely as possible with the characteristics of real-world data.

### v2

@@ 4.1 Identification
$Y_{ui}\left(O_{ui}=1,\bm{O}_{\mathcal{N}_{u}i}=\bm{o}_{\mathcal{N}_{u}i}\right)=Y_{ui}\left(O_{ui}=1,\bm{G}_{ui}=\bm{g}_{ui}\right).$ (3)
Assumption 2
@@
$Y_{ui}\left(O_{ui}=1,\bm{G}_{ui}=\bm{g}_{ui}\right)\perp\left(O_{ui},\bm{G}_{ui}\right)\mid\bm{X}_{ui}.$ (4)
Assumption 3
@@
$\begin{split}&P\left(Y_{ui}\left(O_{ui}=1,\bm{O}_{\mathcal{N}_{u}i}=\bm{o}_{\mathcal{N}_{u}i}\right)=1\mid\bm{X}_{ui}=\bm{x}_{ui}\right)\\ \overset{a}{=}&P\left(Y_{ui}\left(O_{ui}=1,\bm{G}_{ui}=\bm{g}_{ui}\right)=1\mid\bm{X}_{ui}=\bm{x}_{ui}\right)\\ \overset{b}{=}&P\left(Y_{ui}\left(O_{ui}=1,\bm{G}_{ui}=\bm{g}_{ui}\right)=1\mid\bm{X}_{ui}=\bm{x}_{ui},{O}_{ui}=1,\bm{G}_{ui}=\bm{g}_{ui}\right)\\ \overset{c}{=}&P\left(Y_{ui}=1\mid\bm{X}_{ui}=\bm{x}_{ui},{O}_{ui}=1,\bm{G}_{ui}=\bm{g}_{ui}\right).\end{split}$ (6)
Equations (a), (b), and (c) are based on Assumptions 1, 2, and 3, respectively. It is worth noting that the three standard assumptions commonly employed in causal inference are equally applicable in recommendation. Assumption 1 benefits significantly from the strong representational capabilities of graph convolution in recommendation [He2020lightgcn], which can be adapted to capture information regarding the neighbors’ features and the exposure of the item to the neighbors. Assumptions 2 and 3 are adaptations of conventional assumptions in the traditional causal recommendation [wu2022caurec], tailored to account for the presence of interference.
4.2 Interference Representation-based Estimation
Module
@@ 4.2 Interference Representation-based Estimation
$\bm{e}_{ui}=\bm{e}_{u}\odot\bm{e}_{i},$ (7)
where $\bm{e}_{u}$ and $\bm{e}_{i}$ are learnable embeddings for user $u$ and item $i$, respectively, which can be modeled by previous recommender models such as Matrix Factorization (MF) [koren2009mf] and LightGCN [He2020lightgcn].
In social networks, when $u$’s neighbor $u^{\prime}$ is exposed to item $i$, $u^{\prime}$ may be inclined to share its interests $\bm{e}_{u^{\prime}i}$ with $u$. Moreover, the social closeness between different neighbors and a user is not uniform, resulting in differing degrees of influence on the user’s interests in the item. Therefore, the interference representation-based estimation module aggregates unit $(u,i)$’s neighbors’ treatment values $\bm{o}_{\mathcal{N}_{u}i}$ and user $u$’s neighbors’ individual interests in item $i$ as follows:
$\bm{g}_{ui}=\bm{\mathrm{W}}\sum_{u^{\prime}\in\mathcal{N}_{u}}o_{u^{\prime}i}a_{uu^{\prime}}\bm{e}_{u^{\prime}i},$ (8)
@@
$a_{uu^{\prime}}=\frac{\exp(\bm{w}^{\intercal}[\bm{e}_{u}\parallel\bm{e}_{u^{\prime}}])}{\sum_{u^{\prime\prime}\in\mathcal{N}_{u}}\exp(\bm{w}^{\intercal}[\bm{e}_{u}\parallel\bm{e}_{u^{\prime\prime}}])},$ (9)
where $\bm{w}$ is a learnable weight vector and $[\cdot||\cdot]$ denotes the concatenation of two vectors.
Concatenating unit $(u,i)$’s individual representation and interference representation as $\bm{e}_{ui}\parallel\bm{g}_{ui}]$, we can capture user $u$’s final inclination towards interacting with the item $i$. Any classifier can be used to model the interference potential outcome of unit $(u,i)$ (e.g., a linear classifier). In this work, we follow MF and model $P(Y_{ui}(O_{ui}=1,\bm{O}_{\mathcal{N}_{u}i}=\bm{o}_{\mathcal{N}_{u}i})=1\mid\bm{X}_{ui}=\bm{x}_{ui})$ as:
$\sigma\left(\bm{1}^{\intercal}[\bm{e}_{ui}\parallel\bm{g}_{ui}]\right),$ (10)
where $\bm{1}$ is a vector where each element’s value is 1 and $\sigma(\cdot)$ is the sigmoid function. In estimating interference-aware potential outcome, we can specify item $i$’s exposure $\bm{o}_{\mathcal{N}_{u}i}$ to user $u$’s neighbors by observed data $\mathcal{D}$. The modeled interference-aware potential outcomes are denoted as $\hat{p}_{ui}$. We optimize the above learnable parameters by minimizing the following loss:
$\frac{1}{|\mathcal{D}|}\sum_{(\bm{x}_{ui},o_{ui},y_{ui})\in\mathcal{D}}l(\hat{p}_{ui},y_{ui}),$ (11)
@@
4.3 Post-Learning-based Optimization Module
We propose a post-learning-based optimization module to achieve the second point in Section 3.2, which consists of: (1) estimating the sum of probabilities of positive potential outcomes of target units under interference; (2) designing a cost function to measure the damage to the neighbors resulting from the target item’s exposure; and (3) adjusting the treatment values of the neighbors to trade off steering performance and the damage to the neighbors.
4.3.1 Estimation for target units
@@
4.3.2 Design of cost function
We utilize an existing recommender model as the user experience model [^1] In practice, we train a LightGCN [He2020lightgcn], a representative recommender model, as a user experience model, since we currently lack access to the recommendation platforms’ proprietary recommender models designed for user experience. and based on this, we design a cost function. The cost function measures the damage to a neighbor by measuring the decrease in the sum of ratings for items in the adjusted recommendation list from NIRec compared to the sum of ratings in the original recommendation list from the user experience model, both ratings as assigned by the user experience model. In particular, given the user experience model, we obtain: (1) an item list with the top k highest ratings to neighbor $u^{\prime}$; and (2) an exposure treatment value $o^{ex}_{u^{\prime}i^{\ast}}$ indicating whether $i^{\ast}$ is in the top k list ($o^{ex}_{u^{\prime}i^{\ast}}=1$) or not ($o^{ex}_{u^{\prime}i^{\ast}}=0$). From the user experience model’s perspective, if we replace an item in the top k item list, it will damage $u^{\prime}$’s experience. Therefore, we design cost function $c(\cdot)$ as follows:
@@ 4.3.2 Design of cost function
4.3.3 Treatment values optimization
@@
$\begin{split}&\sum_{u^{\ast}\in\mathcal{U^{\ast}}}P(Y_{u^{\ast}i^{\ast}}(\bm{O}_{\mathcal{U}^{\ast}i^{\ast}}=\bm{1},\bm{O}_{\mathcal{N}_{\mathcal{U}^{\ast}}i^{\ast}}=\bm{o}_{\mathcal{N}_{\mathcal{U}^{\ast}}i^{\ast}})=1\mid\bm{X}_{u^{\ast}i^{\ast}}=\bm{x}_{u^{\ast}i^{\ast}})\\ &-\gamma\sum_{u^{\prime}\in\mathcal{N}_{\mathcal{U}^{\ast}}}c\left(o_{u^{\prime}i^{\ast}}\right).\end{split}$ (12)
@@
4.3.4 Scalability
NIRec comprises two modules: the interference representation-based estimation module and the post-learning-based optimization module. The interference representation-based estimation module has a time complexity of $O\left(\left|\mathcal{D}\right|\left(k\,d+d^{2}\right)\right)$, where $\left|\mathcal{D}\right|$ is the size of the training data, $k$ is the average number of social neighbors per user, and $d$ is the embedding dimension. Its computational cost scales linearly with both the dataset size and the density of the social network. Given the precomputed interaction probability of the target users for the target item under the initial exposure, the post-learning-based optimization module has a time complexity of $O\left(k\left(Ft-\frac{t(t-1)}{2}\right)\right)$, where $F$ is the number of the neighbors of the target group, $k$ is the number of target users influenced by each neighbor, and $t$ is the number of iterations required for the optimization objective to converge (worst-case $t=F$, but in practice $t<F$). Consequently, this module’s worst-case complexity is $O\left(F^{2}k\right)$ and its best-case complexity is $O\left(Fk\right)$. If, in reality, the growth rate of social network density is lower than that of the network’s scale, NIRec demonstrates good scalability. In the worst-case scenario, however, as the network scale increases, $\left|\mathcal{D}\right|$, $k$, and $F$ all increase linearly, limiting NIRec’s scalability. Future work will focus on further optimizing NIRec’s scalability.
4.3.5 Discussion
Our framework can be extended to other types of networks where indirect influence might also play a role. The key lies in constructing causal graphs corresponding to the networks that relax interference assumptions. For example, in an item-item similarity network, exposing items related to the target item to the target user will influence the user’s feedback on the target item. The corresponding causal graph will include new causal paths: $\bm{X}_{ui^{\prime}},O_{ui^{\prime}}\rightarrow Y_{ui}$, where $i^{\prime}$ is the neighbor item of item $i$ in the item-item similarity network. We can employ heterogeneous graph neural networks to simultaneously model the interference brought by different networks. We will explore this further in future work.
5 EXPERIMENT SETUP
Table 1: The statistics of Ciao, Epinions, Filmtrust.
@@ 5 EXPERIMENT SETUP
#Interfered User-Item Pairs 179,727 55,029 13,992
Since no offline datasets provide users’ feedback under varying neighbor interference, we propose a framework for user feedback simulation. It is noteworthy that semi-simulation is a viable and commonly practiced experimental method in direct steering [zhu2023influ, bi2024proactive, wang2024incorporate, lian2025itmprec] and for recommendation tasks in interactive evaluation settings [sato2019uplift, gao2024filterbubble, sato2020uplift, gao2023rlrec]. Following the user feedback simulation, we present how to select target users and items for PRSN. Finally, we design evaluation metrics and baselines.
5.1 Real-World Datasets
We use two product review datasets, Ciao and Epinions [tang2012ciao1, tang2012ciao2], as well as a movie review dataset Filmtrust [guo2013film], for simulating users’ feedback under varying neighbor interference, which results in three user feedback simulations. These datasets contain user IDs, item IDs, ratings (from 1 to 5), and a directed social network with trust relations. We use the 10-core setting for Ciao and Epinions and the 5-core setting [^2] The number of items in Filmtrust is on the order of 100, which is relatively small. for Filmtrust. The statistics of the datasets are shown in Table 1.
5.2 User Feedback Simulation
@@
5.2.1 Representation Simulation
We simulate representations for both users and items to generate user feedback under neighbor interference. To this end, we train an MF model using the observed ratings to predict and complete missing ratings in the user-item matrix. However, the model always falls into overestimation on missing ratings [guo2021mrdr]. To this end, we follow [guo2021mrdr] to align the distribution of the ratings in the completed matrix with that of ratings in the random exposure dataset (e.g., Yahoo!R3 [marlin2012yahoo]), thereby enhancing the reliability of our simulations. Thereafter, we use all ratings to train a new MF model to obtain our simulated user representation $\bm{\mathrm{E}}_{\mathcal{U}}^{\ast}$ and item representation $\bm{\mathrm{E}}_{\mathcal{I}}^{\ast}$.
5.2.2 User Feedback Generation Function
According to standard potential outcome simulation in causal inference [ma2021interference, ma2022hypergraph], we design a user feedback generation function considering neighbor interference. Given $\bm{\mathrm{E}}_{\mathcal{U}}^{\ast}$ and $\bm{\mathrm{E}}_{\mathcal{I}}^{\ast}$, we set the positive potential outcome probability of unit $(u,i)$ as:
$\begin{split}\sigma\left(\eta\left(\bm{e}^{\ast}_{u},\bm{e}^{\ast}_{i}\right)+\beta\Delta\left(\bm{\mathrm{E}}^{\ast}_{\mathcal{N}_{u}},\bm{e}^{\ast}_{u},\bm{e}^{\ast}_{i},\bm{o}_{\mathcal{N}_{u}i}\right)\right),\end{split}$ (13)
where $\eta(\bm{e}^{\ast}_{u},\bm{e}^{\ast}_{i})$ is user $u$’s individual interests in interacting with item $i$, which can be shortened to $\eta_{ui}$ for brevity. $\Delta(\bm{\mathrm{E}}^{\ast}_{\mathcal{N}_{u}},\bm{e}^{\ast}_{u},\bm{e}^{\ast}_{i},\bm{o}_{\mathcal{N}_{u}i})$ is the neighbors’ interference effect, which is shortened to $\Delta_{ui}$. $\beta$ is the coefficient that controls the strength of the interference. In particular, we design $\eta_{ui}$ and $\Delta_{ui}$ under a linear setting:
$\begin{split}\eta_{ui}&=\bm{e}^{\ast}_{u}\cdot\bm{e}^{\ast}_{i}-3,\\ \Delta_{ui}&=\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}}\sum_{u^{\prime}\in\mathcal{N}_{u}}{o_{u^{\prime}i}\sigma\left(\bm{e}_{u}^{\ast}\cdot\bm{e}_{u^{\prime}}^{\ast}\right)\eta_{u^{\prime}i}}.\end{split}$ (14)
Here, we consider a user to like an item if the corresponding rating is above 3; otherwise, the user is considered to dislike the item, given that the range of ratings fitted by MF for simulation is from 1 to 5. If the user’s neighbor’s rating of the item is higher than 3, the influence on the user is positive; otherwise, it is negative.
5.2.3 Semi-Synthetic Datasets
We generate synthetic datasets with user feedback to train our interference representation-based estimation module. We adjust the propensity for observing the feedback to ensure that the generated data aligns as closely as possible with the characteristics of real-world data.